# PISA-Bench: The PISA Index as a Multilingual and Multimodal Metric for the Evaluation of Vision-Language Models

### v1

@@ Appendix C Approximating Model PISA Scores
PISA test scores are not reported as solve ratios but are instead measured on the PISA scale using item response theory (IRT) (Hambleton and Swaminathan, 2013; Okubo, 2022). To enable a direct comparison between model performance and student ability on the PISA scale, we estimate an approximate PISA-equivalent score using a Rasch model formulation (Rasch, 1980). We were able to find the official PISA difficulty parameters for a subset of questions in PISA-Bench. We treat the model’s binary correctness outcomes as responses to Rasch-parameterized items and infer the latent ability parameter $\theta$ by maximizing the Bernoulli log-likelihood under the logistic Rasch model. This inferred ability is then mapped to the PISA scale, providing an interpretable, though imperfect, score, as such scores are not widely publicly available. However, it helps us to establish a link between model accuracy on real assessment items and the corresponding human performance level.
@@
To enable a direct comparison between model performance and student ability on the PISA scale, we estimate an equivalent PISA score using a Rasch model formulation. For the subset of benchmark items where official PISA difficulty estimates were available, we treat the model’s binary correctness outcomes as responses to Rasch-parameterized items and compute the latent ability parameter $\theta$
Language en de fr it es ch avg
@@
Human [^2] Scores from 2022: https://worldpopulationreview.com/country-rankings/pisa-scores-by-country (accessed 24.10.2025) 🇺🇸465 🇺🇸504 🇩🇪475 🇩🇪480 🇫🇷474 🇫🇷474 🇮🇹471 🇮🇹482 🇪🇸473 🇪🇸474 🇨🇳552 🇨🇳510 485 487
Table 6: Approximate PISA-equivalent model scores compared with recent national averages (2022) for the United States, Germany, Italy, Spain, France, and China. Due to limited item coverage, these values should be viewed as illustrative only and not as validated student-ability estimates.

### v2

@@ Appendix C Approximating Model PISA Scores
PISA test scores are not reported as solve ratios but are instead measured on the PISA scale using item response theory (IRT) (hambleton2013item; okubo2022theoretical). To enable a direct comparison between model performance and student ability on the PISA scale, we estimate an approximate PISA-equivalent score using a Rasch model formulation (rasch1980probabilistic). We were able to find the official PISA difficulty parameters for a subset of questions in PISA-Bench. We treat the model’s binary correctness outcomes as responses to Rasch-parameterized items and infer the latent ability parameter $\theta$ by maximizing the Bernoulli log-likelihood under the logistic Rasch model. This inferred ability is then mapped to the PISA scale, providing an interpretable, though imperfect, score, as such scores are not widely publicly available. However, it helps us to establish a link between model accuracy on real assessment items and the corresponding human performance level.
@@
To enable a direct comparison between model performance and student ability on the PISA scale, we estimate an equivalent PISA score using a Rasch model formulation. For the subset of benchmark items where official PISA difficulty estimates were available, we treat the model’s binary correctness outcomes as responses to Rasch-parameterized items and compute the latent ability parameter $\theta$.
Language en 🇺🇸 de 🇩🇪 fr 🇫🇷 it 🇮🇹 es 🇪🇸 ch 🇨🇳 Avg.
@@
SmolVLM2-256M-Video-Instruct 387 429 494 373 510 373 510 470 526 440 461 373 481 409
SmolVLM2-500M-Video-Instruct 313 337 387 408 341 281 387 332 408 373 341 440 362 361
SmolVLM2-2.2B-Instruct 461 456 387 332 366 373 387 408 444 440 387 440 405 408
@@
Test Takers
2000 493 504 490 484 517 505 457 487 476 493 - - 487 495
2003 483 495 503 491 511 496 466 476 485 481 - - 490 488
2006 474 - 504 495 496 488 462 469 480 461 - - 483 478
2009 487 500 513 498 497 497 483 486 483 481 600 556 511 503
2012 481 498 514 508 495 505 485 490 484 488 613 570 512 510
2022 465 504 475 480 474 474 471 482 473 474 552 510 485 487
Table 6: Approximate PISA-equivalent model scores compared with national averages for the United States, Germany, Italy, Spain, France, and China from the past years. We compute the PISA scores for LLMs on a subset of questions for which we were able to obtain difficulty parameters. Thus, the resulting LLM scores should be interpreted as approximate indicators rather than validated test scores.